# Novel scaling laws to derive spatially resolved flare and CME parameters from sun-as-a-star observables

### v1

@@ 4.1 $\mathrm{P_{CME}}$ $\mathrm{P_{flare}}$ scaling law: physical insight
The ‘flare-CME-type-II’ phenomenon has two major sub-events, the ‘reconnection-heating’ event in the low corona and the ’CME - particle acceleration’ event extending the impact of the energetic phenomenon to higher corona and IP space leading to various space weather effects. The populations of electrons that are accelerated and the powering mechanism are different in both sub-events, though the evolution of the sub-events are causally and physically connected. So, robust scaling relations that connect the various observable features of the sub-events of a ‘flare- CME- type-II’ phenomenon remained elusive (see, Sec. 1). However, our physics-informed ‘feature engineering’ (Hastie et al., 2009) provide novel metrics, $\mathrm{P_{CME}}$ and $\mathrm{P_{flare}}$, with a significant correlation between them and with other parameters (Eqns. 1– 3). $\mathrm{P_{CME}}$ and $\mathrm{P_{flare}}$ are defined as proxies for the mean power in each sub-event using direct observables: $\mathrm{P_{flare}}$ for the ‘reconnection-heating’ event and $\mathrm{P_{CME}}$ for the ‘CME-particle acceleration’ event. The reconnected magnetic flux ($\mathrm{\phi_{rec}}$) leads to strong local electric fields accelerating electrons in the coronal post-flare loops leading to local heating and associated thermal SXR flares. So $\mathrm{L_{X}}$ is a direct proxy to the heating power. Meanwhile, $\mathrm{\phi_{rec}}$ is a direct observable proxy to the power in the reconnection-driven electric fields. $\mathrm{P_{flare}}$ estimates the coupled power in the reconnection and heating, providing a proxy to the power in the ‘reconnection-heating’ event. Meanwhile, the erupted CME following the reconnection event generates a propagating shock, accelerating electrons and causing the type-II burst. The direct observable proxy to the CME shock power is the $\mathrm{V_{CME}}$2 and $\mathrm{L_{R}}$ is the proxy to the particle acceleration strength. Like $\mathrm{P_{flare}}$, $\mathrm{P_{CME}}$ estimates the typical power of the ‘CME-particle acceleration’ event. The novel power metrics, $\mathrm{P_{flare}}$ and $\mathrm{P_{CME}}$, thus couple the strength of local particle acceleration with the power in their respective drivers, for the two major sub-events of a ‘flare-CME-type-II’ phenomenon.
@@ 4.2 Comparison with GBR
The $\mathrm{P_{CME}}$- $\mathrm{P_{flare}}$ scaling law can be philosophically compared to the GBR. GBR relates the power in the low coronal particle acceleration ($\mathrm{L_{R}}$${}_{\rm,5-8\,GHz}$) and subsequent heating ($\mathrm{L_{X}}$). Any attempted extension or a search for a similar $\mathrm{L_{R}}$-$\mathrm{L_{X}}$ scaling has not been quite successful in the meterwave band (e.g. Callingham et al., 2021; Yiu et al., 2024). A major physical reason for this is that the metric bursts have a variety of dynamic spectral types, with each type having a unique driving mechanism. Here we choose the type-II bursts since they are highly sought after in solar and stellar CME studies. Based on the power in the different drivers of the radio and SXR emission in a ‘flare-CME-type-II’ phenomena, we modify the $\mathrm{L_{R}}$ term in GBR to $\sqrt{\mathrm{L_{R}V^{2}_{CME}}}$ and the $\mathrm{L_{X}}$ to $\sqrt{\mathrm{L_{X}\phi_{rec}}}$. Interestingly, the new power terms show a power law index similar to the GBR, while also accounting for the differing SXR and radio emission formation scenarios (Guedel et al., 1995). A similar careful analysis for other radio burst types may also provide meaningful correlations involving $\mathrm{L_{R}}$ and $\mathrm{L_{X}}$.
@@ 4.3 Implications to stellar CME studies
Being disk-averaged, $\mathrm{L_{X}}$ and $\mathrm{L_{R}}$ are observable in other stars. Combining these observables and the scaling laws associated with the strong correlations (Eqn. 1, and Eqn. 3 or the $\mathrm{L_{X}}$-$\mathrm{\phi_{rec}}$ scaling law estimated over a larger sample (Pevtsov et al., 2003)), the unknown $\mathrm{V_{CME}}$ and $\mathrm{\phi_{rec}}$ can be estimated. $\mathrm{V_{CME}}$ and $\mathrm{\phi_{rec}}$ cannot be directly inferred in stars since flare imaging observations are impossible. Besides, since the type-II emission forms primarily via a plasma emission mechanism, the observation frequency ($\nu$) relates to the local electron density (ne; $\nu\propto\sqrt{\mathrm{n_{e}}}$(Ginzburg & Zhelezniakov, 1958; Melrose, 1970), the type-II frequency drift-rate ($\mathrm{\delta\nu/\delta t}$) and the $\mathrm{V_{CME}}$ estimate can provide the density gradient ($\mathrm{\nabla_{h}n_{e}}$) across the coronal height (h).
@@
where $\mathrm{V_{CME}}$ =$\mathrm{\delta h/\delta t}$. $\mathrm{V_{CME}}$, $\mathrm{\phi_{rec}}$  and $\mathrm{\nabla_{h}n_{e}}$ are crucial constraints to flare and CME evolution models.
Since DH type-IIs and metric bursts have similar emission mechanism, the results presented provide a direction to extend the $\mathrm{L_{R}}$- $\mathrm{L_{X}}$ correlation studies to the metric band by carefully considering the physical drivers of SXR and radio emission. The extension of the analysis presented here to metric waveband is beyond the scope of this work given the lack of uniformly calibrated long term metric dynamic spectral database, unlike the DH band. However, the results presented provide a scientific motivation to calibrate and explore the metric type-II events using a similar methodology. But, one may have to use the CME speed in the lower coronal heights probed by the meterwave band as opposed to the IP space $\mathrm{V_{CME}}$ used here.
Given the low occurrence rate of CME-associated metric bursts in active stars (Villadsen & Hallinan, 2019; Zic et al., 2020; Mohan et al., 2024b), and the expected higher occurrence rate in the DH band (Alvarado-Gómez et al., 2022), there has been a push for space-based DH band observatories. Our results aid this initiative by providing the flux estimates of the ‘flare-CME-type-II’ events with varying $\mathrm{V_{CME}}$- $\mathrm{\phi_{rec}}$ values.
@@ 5 Conclusion
We present a novel scaling law for ‘flare- CME- type-II’ events using decades-long solar decameter-hectrometric type-II burst data and simultaneous multi-waveband imaging and non-imaging database. The scaling relation links two novel metrics of ‘flare power’ ($\mathrm{P_{flare}}$= $\sqrt{\mathrm{L_{X}\phi_{rec}}}$) and ‘CME power’ ($\mathrm{P_{CME}}$= $\sqrt{\mathrm{L_{R}V^{2}_{CME}}}$), where $\mathrm{L_{X}}$ is the SXR luminosity, $\mathrm{L_{R}}$ the type-II luminosity, $\mathrm{\phi_{rec}}$ is the reconnection flux and $\mathrm{V_{CME}}$ is the CME speed. When $\mathrm{P_{flare}}$ encapsulates the power in low-coronal heating and the associated SXR flare, $\mathrm{P_{CME}}$ relates to CME shock-driven particle acceleration and the associated type-II burst. The two terms together encapsulate the two major manifestations of ‘flare- CME- type-II’ event in the low-coronal (‘reconnection-heating’) and interplanetary (‘CME-particle acceleration’) regions. $\mathrm{P_{flare}}$$\propto$ $\mathrm{P_{CME}}$0.76. Analogous to the Güdel Benz relation (GBR) that links the flare thermal power and the particle acceleration strength driven by the low-coronal reconnection event, the novel law derives a connection between the low coronal and interplanetary (IP) particle acceleration driven by the reconnection and the CME events. Previous attempts to identify GBR-like scaling laws connecting $\mathrm{L_{X}}$ and $\mathrm{L_{R}}$ in metric band using stellar data had not been successful. This study provides a philosophy and approach to extend an $\mathrm{L_{X}}$-$\mathrm{L_{R}}$ relationship for long-wavelengths by taking into account the emission drivers. Additionally, we derive scaling laws linking $\mathrm{L_{X}}$ and $\mathrm{\phi_{rec}}$ to $\mathrm{P_{CME}}$. By combining the derived and known scaling laws with disk-integrated $\mathrm{L_{X}}$ and $\mathrm{L_{R}}$ observable in stellar events, $\mathrm{V_{CME}}$, $\mathrm{\phi_{rec}}$ can be estimated. Additionally assuming a plasma emission mechanism for type-II bursts, the frequency drift rate can be used to estimate $\mathrm{\nabla_{h}n_{e}}$ in the corona, once $\mathrm{V_{CME}}$ is estimated. The presented study takes a top-down approach to understand the interconnections between the two major sub-events of the ‘flare-CME-type-II’ phenomena. There need to be a bottom-up approach using modeling to understand the emergence of the various scaling laws, particularly the $\mathrm{P_{flare}}$-$\mathrm{P_{CME}}$ scaling.
@@ Acknowledgements.
AM, NG and HR are partly supported by the NASA’s STEREO project and LWS program. SA was partially supported by NSF grant, AGS-2043131. We thank the CDAW team for maintaining an up-to-date catalog [^3] https://cdaw.gsfc.nasa.gov/CME_list/ of solar CMEs detected by the Large Angle and Spectrometric Coronagraph (LASCO) on board the Solar and Heliospheric Observatory (SOHO) mission. AM acknowledges Pertti Mäkelä for maintaining an up-to-date DH type-II catalog. AM acknowledges Vratislav Krupar and NASA’s Space Physics Data Facility (SPDF) [^4] https://spdf.gsfc.nasa.gov/ for the calibrated radio dynamic spectra. AM acknowledges the developers of the various Python modules namely Numpy (Harris et al., 2020), Astropy (Astropy Collaboration et al., 2013), Matplotlib (Hunter, 2007) and Pandas (pandas development team, 2020).

### v2

@@ 4.1 $\mathrm{P_{CME}}$ $\mathrm{P_{flare}}$ scaling law: Physical insight
The flare-CME-type II phenomenon has two major sub-events, the ”reconnection-heating” event in the low corona and the ”CME - particle acceleration” event extending the impact of the energetic phenomenon to higher corona and IP space leading to various space weather effects. The populations of accelerated electrons and the powering mechanism are different in both sub-events, though the evolution of the sub-events is causally and physically connected. As a result, robust scaling relations that connect the various observable features of the sub-events of a flare-CME-type II phenomenon remained elusive (see, Sec. 1). However, our physics-informed ”feature engineering” (Hastie et al., 2009) provides novel metrics, $\mathrm{P_{CME}}$ and $\mathrm{P_{flare}}$, with a significant correlation between them and with other parameters (Eqns. 1– 3). The $\mathrm{P_{CME}}$ and $\mathrm{P_{flare}}$ are defined as proxies for the mean power in each sub-event using direct observables: $\mathrm{P_{flare}}$ for the reconnection-heating event and $\mathrm{P_{CME}}$ for the CME-particle acceleration event. The reconnected magnetic flux ($\mathrm{\phi_{rec}}$) leads to strong local electric fields accelerating electrons in the coronal post-flare loops leading to local heating and associated thermal SXR flares. So $\mathrm{L_{X}}$ is a direct proxy for the heating power. Meanwhile, $\mathrm{\phi_{rec}}$ is a direct observable proxy for the power in the reconnection-driven electric fields. We note that $\mathrm{P_{flare}}$ estimates the coupled power in the reconnection and heating, providing a proxy for the power in the reconnection-heating event. Meanwhile, the erupted CME following the reconnection event generates a propagating shock, accelerating electrons and causing the type II burst. The direct observable proxy for the CME shock power is the $\mathrm{V_{CME}}$2 and $\mathrm{L_{R}}$ is the proxy for the particle acceleration strength. Similar to $\mathrm{P_{flare}}$, $\mathrm{P_{CME}}$ estimates the typical power of the CME-particle acceleration event. The novel power metrics, $\mathrm{P_{flare}}$ and $\mathrm{P_{CME}}$, thus couple the strength of local particle acceleration with the power in their respective drivers for the two major sub-events of a flare-CME-type II phenomenon.
@@ 4.2 Comparison with GBR
The $\mathrm{P_{CME}}$- $\mathrm{P_{flare}}$ scaling law can be philosophically compared to the GBR. GBR relates the power in the low coronal particle acceleration ($\mathrm{L_{R}}$${}_{\rm,5-8\,GHz}$) and subsequent heating ($\mathrm{L_{X}}$). Any attempted extension or a search for a similar $\mathrm{L_{R}}$-$\mathrm{L_{X}}$ scaling has not quite been successful in the meterwave band (e.g., Callingham et al., 2021; Yiu et al., 2024). A major physical reason for this is that the metric bursts have a variety of dynamic spectral types, with each type having a unique driving mechanism. For this work we chose the type II bursts since they are highly sought after in solar and stellar CME studies. Based on the power in the different drivers of the radio and SXR emission in flare-CME-type II phenomena, we modified the $\mathrm{L_{R}}$ term in GBR to $\sqrt{\mathrm{L_{R}V^{2}_{CME}}}$ and the $\mathrm{L_{X}}$ to $\sqrt{\mathrm{L_{X}\phi_{rec}}}$. Interestingly, the new power terms show a power-law index similar to the GBR, while also accounting for the differing SXR and radio emission formation scenarios (Guedel et al., 1995). A similar careful analysis for other radio burst types may also provide meaningful correlations involving $\mathrm{L_{R}}$ and $\mathrm{L_{X}}$.
@@ 4.3 Implications to stellar CME studies
Being disk-averaged, $\mathrm{L_{X}}$ and $\mathrm{L_{R}}$ are observable in other stars. Combining these observables and the scaling laws associated with the strong correlations (Eqn. 1, and Eqn. 3 or the $\mathrm{L_{X}}$-$\mathrm{\phi_{rec}}$ scaling law estimated over a larger sample (Pevtsov et al., 2003)), the unknown $\mathrm{V_{CME}}$ and $\mathrm{\phi_{rec}}$ can be estimated. We note that $\mathrm{V_{CME}}$ and $\mathrm{\phi_{rec}}$ cannot be directly inferred in stars since flare imaging observations are impossible. In addition, since the type II emission forms primarily via a plasma emission mechanism, the observation frequency ($\nu$) relates to the local electron density (ne; $\nu\propto\sqrt{\mathrm{n_{e}}}$(Ginzburg & Zhelezniakov, 1958; Melrose, 1970), the type II frequency drift rate ($\mathrm{\delta\nu/\delta t}$) and the $\mathrm{V_{CME}}$ estimate can provide the density gradient ($\mathrm{\nabla_{h}n_{e}}$) across the coronal height (h):
@@
where $\mathrm{V_{CME}}$ =$\mathrm{\delta h/\delta t}$. Furthermore, $\mathrm{V_{CME}}$,  $\mathrm{\phi_{rec}}$ , and $\mathrm{\nabla_{h}n_{e}}$ are crucial constraints to flare and CME evolution models.
Since DH type IIs and metric bursts have a similar emission mechanism, the results presented provide a direction to extend the $\mathrm{L_{R}}$- $\mathrm{L_{X}}$ correlation studies to the metric band by carefully considering the physical drivers of SXR and radio emission. The extension of the analysis presented here to the metric waveband is beyond the scope of this work given the lack of a uniformly calibrated long-term metric dynamic spectral database, unlike the DH band. However, the results presented provide a scientific motivation to calibrate and explore the metric type II events using a similar methodology. However, one may have to use the CME speed in the lower coronal heights probed by the meterwave band as opposed to the IP space $\mathrm{V_{CME}}$ used here.
Given the low occurrence rate of CME-associated metric bursts in active stars (Villadsen & Hallinan, 2019; Zic et al., 2020; Mohan et al., 2024b), and the expected higher occurrence rate in the DH band (Alvarado-Gómez et al., 2022), there has been a push for space-based DH band observatories. Our results aid this initiative by providing the flux estimates of the flare-CME-type II events with varying $\mathrm{V_{CME}}$- $\mathrm{\phi_{rec}}$ values.
@@ 5 Conclusion
We present a novel scaling law for flare-CME-type II events using decades-long solar decameter-hectometric type II burst data and the simultaneous multi-waveband imaging and non-imaging database. The scaling relation links two novel metrics of flare power ($\mathrm{P_{flare}}$= $\sqrt{\mathrm{L_{X}\phi_{rec}}}$) and CME power ($\mathrm{P_{CME}}$= $\sqrt{\mathrm{L_{R}V^{2}_{CME}}}$), where $\mathrm{L_{X}}$ is the SXR luminosity, $\mathrm{L_{R}}$ is the type II luminosity, $\mathrm{\phi_{rec}}$ is the reconnection flux, and $\mathrm{V_{CME}}$ is the CME speed. When $\mathrm{P_{flare}}$ encapsulates the power in low-coronal heating and the associated SXR flare, $\mathrm{P_{CME}}$ relates to CME shock-driven particle acceleration and the associated type II burst. The two terms together encapsulate the two major manifestations of the flare-CME-type II event in the low-coronal (reconnection-heating) and interplanetary (CME-particle acceleration) regions. We note that $\mathrm{P_{flare}}$$\propto$ $\mathrm{P_{CME}}$0.76. Analogous to the GBR, which links the flare thermal power and the particle acceleration strength driven by the low-coronal reconnection event, the novel law derives a connection between the low coronal and IP particle acceleration driven by the reconnection and the CME events. Previous attempts to identify GBR-like scaling laws connecting $\mathrm{L_{X}}$ and $\mathrm{L_{R}}$ in the metric band using stellar data had not been successful. This study provides a philosophy and approach to extend an $\mathrm{L_{X}}$-$\mathrm{L_{R}}$ relationship for long wavelengths by taking the emission drivers into account. Additionally, we derived scaling laws linking $\mathrm{L_{X}}$ and $\mathrm{\phi_{rec}}$ to $\mathrm{P_{CME}}$. By combining the derived and known scaling laws with disk-integrated $\mathrm{L_{X}}$ and $\mathrm{L_{R}}$ observable in stellar events, $\mathrm{V_{CME}}$  and $\mathrm{\phi_{rec}}$ can be estimated. Additionally, assuming a plasma emission mechanism for type II bursts, the frequency drift rate can be used to estimate $\mathrm{\nabla_{h}n_{e}}$ in the corona, once $\mathrm{V_{CME}}$ is estimated. The presented study takes a top-down approach to understanding the interconnections between the two major sub-events of the flare-CME-type II phenomena. There needs to be a bottom-up approach using modeling to understand the emergence of the various scaling laws, particularly the $\mathrm{P_{flare}}$-$\mathrm{P_{CME}}$ scaling.
@@ Acknowledgements.
AM, NG and HR are partly supported by NASA’s STEREO project and LWS program. SA was partially supported by NSF grant, AGS-2043131. We thank the CDAW team for maintaining an up-to-date catalog [^4] https://cdaw.gsfc.nasa.gov/CME_list/ of solar CMEs detected by the Large Angle and Spectrometric Coronagraph (LASCO) on board the Solar and Heliospheric Observatory (SOHO) mission. AM acknowledges Pertti Mäkelä for maintaining an up-to-date DH type II catalog. AM acknowledges Vratislav Krupar and NASA’s Space Physics Data Facility (SPDF) [^5] https://spdf.gsfc.nasa.gov/ for the calibrated radio dynamic spectra. AM acknowledges the developers of the various Python modules namely Numpy (Harris et al., 2020), Astropy (Astropy Collaboration et al., 2013), Matplotlib (Hunter, 2007), and Pandas (pandas development team, 2020).